OpenStudy (anonymous):

Need help with implicit differentiation! xy^2+yx^2=6

OpenStudy (anonymous):

What do you mean by partial derivatives? Should I use the product rule?

OpenStudy (anonymous):

this is an implicit differentiation

OpenStudy (anonymous):

But this requires implicit differentiation, so you rather treat y as a function of x. I recommend to rewrite the equation this way.

OpenStudy (anonymous):

so u need the product rule as well

OpenStudy (anonymous):

\[\Large x y(x)^2+y(x)x^2=6 \]

OpenStudy (anonymous):

Where do I take it from there?

OpenStudy (anonymous):

\[2xy \frac{dy}{dx}+y^2+\frac{dy}{dx}x^2+2xy=0\]

OpenStudy (anonymous):

that is what we call implicit differentiation...

OpenStudy (anonymous):

this is not partial diff...

OpenStudy (anonymous):

Can you please explain why you pu in the dy/dx?

OpenStudy (anonymous):

because the question says find the implicit differentiation, which a fondamental subject for differential equations... now if you mean how did i get that result, well that is simple to see,,,

OpenStudy (anonymous):

Ok I have \[P _{0}\] (1,2) also. Where do I plug those in?

OpenStudy (anonymous):

first of all xy^2+yx^2=6 as you can see x is multiplied by y^2 so you need to use the product rule, similarly you have yx^2 which requires product rule of differentiation...

OpenStudy (anonymous):

next step you need to seperate variables...

OpenStudy (anonymous):

what is Po? you need to provide more information, is it the tangent point, what is it?

OpenStudy (anonymous):

Can I add 2xy\[\frac{ dy }{ dx }\] + 2xy or is thattwo completely different things. The question says: Use the method of implicit differentiation to calculate dy/dx at the point Po

OpenStudy (anonymous):

ok so it is the slope of the tangent, so now you need to solve for dy/dx, and yes you can add 2xy... like solving a normal algabraec problem.,..

OpenStudy (anonymous):

that mean dy/dx should = something...

OpenStudy (anonymous):

if you need help let me know, put dy/dx on one side and the rest on the other side...

OpenStudy (anonymous):

or just plug in the numbers as it is, that would be easier for you

OpenStudy (anonymous):

put for every x=1 and for every y=2

OpenStudy (anonymous):

don't waste your time on rearranging the formula...

OpenStudy (anonymous):

\[2(1)(2) \frac{dy}{dx}+(2)^2+\frac{dy}{dx}(1)^2+2(1)(2)=0\]

OpenStudy (anonymous):

Okay, when I plug in the x and y to the formula, with out rearranging it, i don't plug it into dy/dx also do I?

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

ok so now i solve and leave the dy/dx in?

OpenStudy (anonymous):

yes dy/dx is what you want to find at that point...

OpenStudy (anonymous):

ok so after i solved I end up with 4 dy/dx + dy/dx +8. How do I find dy/dx?

OpenStudy (anonymous):

take dy/dx as your common factor

OpenStudy (anonymous):

Sorry I still don't understand

OpenStudy (anonymous):

Oh so 5 dy/dx + 8 =0?

OpenStudy (anonymous):

It worked!! Thank you SO MUCH!

OpenStudy (anonymous):

\[4\frac{dy}{dx}+\frac{dy}{dx}+8=0 \rightarrow 5\frac{dy}{dx}=-8\]

OpenStudy (anonymous):

mr wislar stop those pictures please

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Ok so at the beginning of the problem, how do I know when to incorporate in the dx/dy?

OpenStudy (anonymous):

your final answer is dy/dx=-8/5

OpenStudy (anonymous):

Yes I got that! Thanks! But now i'm moving on and I don't understand how you got the equation 2xy dy/dx + y^2 + dy/dx x^2

OpenStudy (anonymous):

+2xy

OpenStudy (anonymous):

ok i will show you

OpenStudy (anonymous):

consider the following example i want to avoid using termonologies ok...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x^4+y^4=4\]

OpenStudy (anonymous):

\[\frac{d(x^4)}{dx} + \frac{d(y^4)}{dx}=\frac{d(4)}{dx}\]

OpenStudy (anonymous):

4x^3 is the first part do you agree with that?

OpenStudy (anonymous):

diff x^4 =4x^3

OpenStudy (anonymous):

diff(4)=0 do you agree?

OpenStudy (anonymous):

now when you want to differentiat y^4 with respect to x, you get into the chain rule world

OpenStudy (anonymous):

diff(y^4)=(4y^3)y'

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the only new thing is when you differentiate the y?

OpenStudy (anonymous):

we will report an abuse for any distraction ,,, stop sending pictures

OpenStudy (anonymous):

we don't want viruses

OpenStudy (anonymous):

Why did you multiply dy/dx to y, but add dy/dx to x?

OpenStudy (anonymous):

sorry I mean y^2 and x^2

OpenStudy (anonymous):

ok i will prove that for you it is based on the chain rule principle combined with the product rule...

OpenStudy (anonymous):

the prodcut rule says you differentiate the first part leave the 2nd function , then it says leave the first part and differentiate the 2nd function

OpenStudy (anonymous):

\[D(x^2y^2)= x^2 2yy'+2y^2x\]

OpenStudy (anonymous):

did u get that?

OpenStudy (anonymous):

Ok I think so

OpenStudy (anonymous):

ok in a normal case how would diff xe^x?

OpenStudy (anonymous):

Would we use the chain rule?

OpenStudy (anonymous):

no the product rule...

OpenStudy (anonymous):

when you have two functions muliplied by each other you use the product rule

OpenStudy (anonymous):

Thank you so much for your patience and help! I have to go because my battery is dying. Your help is greatly appreciated and I did learn from you thank you so much again!

OpenStudy (anonymous):

you're more than welcome! all the best

OpenStudy (anonymous):